Parth (parthkohli):

How do I prove that if the sum of two numbers is \(x\), then \(\frac{x}{2}, \frac{x}{2}\) maximizes the product?

OpenStudy (anonymous):

Is it x/2 , x/2 ?

Parth (parthkohli):

Yes

Parth (parthkohli):

Let the two numbers be \((x - n)(x -k) = x^2 - nx - kx + nk= x^2 - (n+k) x + nk\). I don't know if that helps.

OpenStudy (anonymous):

square rule helps

Parth (parthkohli):

I just have to maximize \(x^2 - (n + k)x + nk\) but have no idea how.

OpenStudy (zarkon):

a+b=x a*b=max use calculus

OpenStudy (anonymous):

i wouldn't

OpenStudy (anonymous):

i would argue by symmetry call one \(a\) and the other \(b\) with \(a+b=x\) then since this equation is symmetric in \(a\) and \(b\), i.e. you cannot tell them apart, it must be largest in the middle

Parth (parthkohli):

But it can also be the smallest

OpenStudy (anonymous):

by which i mean the product must be largest in the middle but you can write as @Zarkon said maximize \(a(x-a)\) using calculus

OpenStudy (anonymous):

can't be the smallest if \(a=x\) or \(b=x\) you get \(ab=0\)

OpenStudy (anonymous):

i.e. it is smallest at the endpoint of the interval

OpenStudy (anonymous):

or for that matter you can find the max of \(ax-a^2\) is at the vertex, which gives \(\frac{x}{2}\)

Parth (parthkohli):

Yeah, that's a good idea.

OpenStudy (anonymous):

Let those numbers be a and b a + b = x You have to prove that : \(ab < \cfrac{x^2}{4} \) a + b = x \(\cfrac{a+b}{2} = \cfrac{x}{2}\) Using A.M - G.M inequality : \(\cfrac{\alpha + \beta}{2} > \sqrt{\alpha \beta}\) Therefore : \(\cfrac{a+b}{2} > \sqrt{ab} \) \(\cfrac{x}{2} > \sqrt{ab}\) \(\cfrac{x^2 }{4} > ab\) or : \(ab < \cfrac{x^2}{4} \)

OpenStudy (anonymous):

Hence Proved (if that was to prove basically :) )

Parth (parthkohli):

It should be \(\le\) and \(\ge\), but I get your point. Real good :-)

OpenStudy (anonymous):

nah, i like symmetry more look, you call one \(a\) and the other \(b\) where \(a+b=x\) but i come along and call the first one \(b\) and the second one \(a\) and write \(b+a=x\) it is pretty clear that there is no difference between them so how could you favour \(a\) over \(b\) somehow, so that the answer would not be \(a=b\) ?

OpenStudy (anonymous):

Symmetry is good too. But I think using AM - GM inequality is also fine and good. :)

OpenStudy (zarkon):

what if x<0

OpenStudy (anonymous):

then i am probably wrong

OpenStudy (zarkon):

I'm referring to the AM - GM inequality

OpenStudy (anonymous):

actually i think i am still right

OpenStudy (zarkon):

you are

OpenStudy (anonymous):

happens about once a day, sometimes twice

OpenStudy (zarkon):

lol

Parth (parthkohli):

lol

OpenStudy (anonymous):

Done (in a clever 5th grade way) ;)

OpenStudy (anonymous):

Have a look at my solution :

OpenStudy (anonymous):

No problem then : \(\bf { a+b = x\\ \text{Now} \space , (a-b)^2 \ge 0 \\ a^2 + b^2 - 2ab \ge 0 \\ a^2 + b^2 \ge 2ab \\ a^2 + b^2 + 2ab \ge 4ab \\ (a+b)^2 \ge 4ab \\ \cfrac{(a+b)^2}{4} \ge ab \\ \cfrac{x^2}{4} \ge ab \\ } \) \(\textbf{Proved }\) ....

OpenStudy (anonymous):

Looks better now. Isn't it @ParthKohli @Zarkon @satellite73 :)

Parth (parthkohli):

Ah! Great one again!

OpenStudy (anonymous):

Good to hear.